Two point charges are fixed on the y axis: a negative point charge q1 = -29  ?C at y1 = +0.19  m and a positive point charge q2 at y2 = +0.32 m. A third point charge q = +8.4  ?C is fixed at the origin. The net electrostatic force exerted on the charge q by the other two charges has a magnitude of 29  N and points in the +y direction. Determine the magnitude of q2.
Answer #1

I have solved this question earlier with different figures. Please workout using yours figures. If you need any further help just PM me. If I have helped you please rate me 5 stars first (before you rate anyone else).


I have solved this question earlier with different figures. Please workout using yours figures. If you need any further help just PM me. If I have helped you please rate me 5 stars first (before you rate anyone else).


net force=26N= Kq(q2/y2^ 2 + q1/y1^2 );

q2 =(26+ K.q.q1/y1^2).(y2^2/Kq)

=(26 + 9*8.4*26*10^-3 / 0.19^2).(0.31^2 / 9*8.4*10^3)

=1.023*10^-4

=0.1023 mC   
Answer #2

net force=26N= Kq(q2/y2^ 2 + q1/y1^2 );

q2 =(26+ K.q.q1/y1^2).(y2^2/Kq)

=(26 + 9*8.4*26*10^-3 / 0.19^2).(0.31^2 / 9*8.4*10^3)

=1.023*10^-4

=0.1023 mC   
Answer #3

net force=29N= Kq(q2/y2^2 + q1/y1^2 );

=> 29 = 9*10^9 * 8.4*10^-12 [29/0.19^2 - q2/0.32^2]

=> q2 = +42.97 micro C
Answer #4

net force= 29N= Kq(q1/y1^2 + q2/y2^2 );

here k = 9*10^9

q = 8.4*10^-6 C

q1 = -29*10^-6 C

so


=> 29 = 9*10^9 * 8.4*10^-6 [29*10^-6 /0.19^2 - q2/0.32^2]


=> q2 = +42.97 * 10^-6 C

=> q2 = +42.97 micro C